solve undetermined coefficient y''+y=2x+3ex

1
Expert's answer
2021-06-17T09:48:09-0400

"y'' +y=2x+3e^x"

1)"y''+y=0"

"k^2 +1=0"

"k_1 = i, k_2=-i"

"y=C_1\cos(x)+C_2\sin(x)"

"k_1+ik_2=i+1\not= k_1\not= k_2"

"y=e^{k_1}A_n(x)=Ae^{i}"

2)"y''+y=2x+3e^x"

"y=Ae^x+Bx"

"y'= Ae^x+B"

"y'' = Ae^x"

"y'' + y' = Ae^x+Ae^x+Bx=2x+3e^x"

"\begin{cases}
2A=3=>A=3/2\\
B=2
\end{cases}"

"y=C_1\cos(x) +C_2\sin(x)+\dfrac{3}{2}e^x+2x"

Checking:

"y'' = -C_1\cos(x) - C_2\sin(x)+\dfrac{3}{2}e^x"

"y'' + y = -C_1\cos(x) - C_2\sin(x)+\dfrac{3}{2}e^x+C_1\cos(x) +C_2\sin(x)+\dfrac{3}{2}e^x+2x=3e^x+2x"

Answer:

"y=C_1\cos(x) +C_2\sin(x)+\dfrac{3}{2}e^x+2x"